OpenStudy (anonymous):

Factor completely 3x4 – 30x3 + 75x2 3(x – 5)2 3x2(x – 5)2 3x2(x2 – 10x + 25) 3x2(x + 5)(x – 5)

OpenStudy (anonymous):

the 2's are exponents fyi

OpenStudy (cwrw238):

you take the GCF out of brackets : = 3x^2

OpenStudy (cwrw238):

whats left after dividing by GCF = x^2 – 10x + 25 but can this be factored ?

OpenStudy (anonymous):

idk O.o

OpenStudy (cwrw238):

x^2 – 10x + 25 - to factor this we need 2 numbers a and b where ab = + 25 and a + b = -10 these are (same number in this case) -5 and -5 so we get (x - 5)(x - 5) = (x - 5)^2

OpenStudy (anonymous):

thanks for clear explanation!

OpenStudy (cwrw238):

so which choice is the correct one?

OpenStudy (cwrw238):

B right?

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

3x²(x - 5)²

OpenStudy (anonymous):

answer^

OpenStudy (cwrw238):

yes